A doctor prescribes an ointment that is 4% hydrocortisone. A pharmacist has 1% and 5% concentrations in stock. How many ounces of each should the pharmacist use to make a 1 ounce tube?
Answer by mananth(16946)

percent--------- percentage ----------------quantity
Hydrocortisone type I 1---------------x ounce
Hydrocortisone type II 5----------------1-x ounce
Mixture 4.00% ---------------- 1 ounce

1x+5(1- x)=4*1
1x+5-5x =4
x-5x= 4 - 5
-4x=-1
/-4
x=0.25 ounce 1.00% Hydrocortisone type I
0.75 ounce 5.00% Hydrocortisone type II